The infant mortality rate (per 1,000 live births) for the 50 states and the District of Columbia has a mean of 6.98 and a standard deviation of 1.62. Assuming that the distribution is normal, what percentage of states have an infant mortality rate between 5.36 and 8.60 percent (per 1,000 live births)?
Answer #1

answer)

As the given data is normally distributed

We can use standard normal z table to estimate the answer

Given mean = 6.98

S.d = 1.62

Z = (x-mean)/s.d

We need to find

P(5.36<x<8.60) = p(x<8.6) - p(x<5.36)

P(x<8.6)

Z = (8.6 - 6.98)/1.62

Z = 1

From z table, p(z<1) = 0.8413

P(x<5.36)

Z = (5.36 - 6.98)/1.62

Z = -1

From z table

P(z<-1) = 0.1587

Required proportion is

0.8413 - 0.1587

= 0.6826

That is 68.26%

Or approximately 68%